Can you please help me write this as an equation?
one number is 3 more then twice another number.find the larger number if their sum is -15
I think it is written like this;
2x+3=x

Answer by ptaylor(2198)

Let x=one of the numbers
Then 2x+3 = the other number
And we are told that:
2x+3+x=-15
3x=-15-3
3x=-18
x=-6------Larger number
2x+3=2(-6)+3=-12+3=-9-----Smaller number
Hope this helps---ptaylor